OpenStudy (anonymous):

Use basic identities to simplify the expre2ssion. sin^2theta+tan^2theta+cos^2theta

OpenStudy (xapproachesinfinity):

remember that \[\sin^2\theta+\cos^2\theta=1\]

OpenStudy (xapproachesinfinity):

also \[1+\tan^2\theta=\sec^2\theta\]

OpenStudy (xapproachesinfinity):

figure out the rest...

OpenStudy (jdoe0001):

ditto

OpenStudy (welshfella):

yea

OpenStudy (jdoe0001):

as easy as \(\huge \pi\)